40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 1 22:56:27 2014, Koji, Update, General, Beam lost in the chamber??? 
    Reply  Sun Aug 3 17:19:32 2014, ericq, Update, General, Recovery efforts 
       Reply  Sun Aug 3 23:47:56 2014, Koji, Update, General, Recovery efforts 
          Reply  Mon Aug 4 17:30:00 2014, ericq, Update, General, Chronic Suspension Problems 
             Reply  Mon Aug 4 18:25:46 2014, Jenne, Update, General, Chronic Suspension Problems 
    Reply  Mon Aug 4 09:17:42 2014, Steve, Update, SUS, ITMX is still kicked ITMXkicked.pngITMXkickedStart.png
Message ID: 10326     Entry time: Sun Aug 3 17:19:32 2014     In reply to: 10325     Reply to this: 10327
Author: ericq 
Type: Update 
Category: General 
Subject: Recovery efforts 

[ericq, Jenne]

We both happened to come by today to fix things up.

When I arrived, the PMC was locked to a 01 mode, which I fixed. The PMC transmission is still worryingly low. MC locked happily. 

ETMX was getting odd kicks, the kind where a DC shift would occur suddenly, and then go away a few moments later. I turned off all dynamic coil outputs, and looked at the MON output of the SOS driver with a scope to try and see if the DAC or dewhitening was glitching, but didn't see anything... Meanwhile, Jenne fiddled with the TTs until we got beams on POP and REFL. (EDIT, JCD:  Useful strategies were to put an excitation onto TT2, and move TT1 until the scattered beam in the chamber was moving at the excitation frequency,  Find the edges of TT2 by finding where the scattered light stops seeing the excitation, and center the beam on TT2.  By then, I think I saw the beam on the PRM face camera.  Then, put a temporary camera looking at the face of PR2.  Using TT2 to center here got us the beam on the POP camera.)

We then walked PRM and the TTs around to keep those two camera beams and get the PRM oplev beam back on its QPD. At this point, ITMX was misaligned (by us), and ITMY aligned to get some recycled flashes into the Y-arm. Y-arm was locked to green, and we poked TTs to get better IR flashes. Misaligning PRM, we had Y-Arm flashes of ~0.7. From there, the michelson and then X-arm were roughly aligned. Both arms were seeing flashes of about 0.7, and the MICH fringes on the AS port look nice.

Frustratingly, the SUS->LSC communication for TRY and TRX isn't working, and could not be fixed by any combination of model or front-end restarting... Thus we haven't been able to actually lock the arms and run ASS. THIS IS VERY FRUSTRATING. 

Additionally, at the point where we were getting light back into the Yarm, the ITMX that were seen on Friday were happening again, tripping the watchdog. Also, something in the Yarm cavity is getting intermittently pushed around, as can be seen by the green lock suddenly wandering off. All of these suspension shenanigans seem to be independent of oplev damping. 

It troubles me that this whole situation is fairly similar to the last time we lost the input pointing (ELOG 10088)

In any case, we feel that we have gotten the IFO alignment to a lockable state.

ELOG V3.1.3-